# A Hybrid Machine Learning Approach for Synthetic Data Generation with Post Hoc Calibration for Clinical Tabular Datasets

### v1

@@ Abstract
Healthcare AI, Synthetic Data Generation, Privacy Preserving Techniques, Machine Learning Models, Data Augmentation, Calibration Methods
@@ 1 Introduction
Data-driven healthcare depends on access to high-quality, representative datasets for building and validating robust AI models [1]. Privacy regulations in the United States impose strict controls on health data use and sharing [2]. European regulations further constrain collection, processing, and secondary use of personal data, including medical records [3]. Synthetic data generation offers a practical pathway by producing artificial datasets that mirror the statistical properties of real data while reducing disclosure risk [4]. In this study, we analyze the Breast Cancer Wisconsin (Original) and Breast Cancer Wisconsin (Diagnostic) datasets from the UCI Machine Learning Repository and a heart-disease cohort from the Department of Cardiology, Khulna Medical College [5].
Synthetic data enables applications that would otherwise be limited by access barriers in clinical research [6]. It can support evaluation of treatment strategies and simulation of clinical workflows without exposing patient identities [7]. Choi et al. developed medGAN to synthesize multi-label discrete electronic health records for predictive modeling when real-world data are scarce [8]. In medical imaging, Frid-Adar et al. showed that GAN-based augmentation can improve diagnostic performance in limited-sample settings [9]. These developments illustrate how synthetic data can advance AI while adhering to ethical and legal expectations [10].
@@
Despite this promise, generating high-fidelity tabular data that respects complex relationships remains challenging [11]. Classical probabilistic models can struggle with high dimensionality and nonlinear structure [12]. GAN-based approaches have improved realism for many data types [13]. Variational autoencoders provide a complementary probabilistic formulation for synthesis [14]. Nonetheless, issues such as mode collapse, overfitting, and data scarcity can hinder practical deployment in sensitive domains [15]. Careful post-generation calibration is often required to preserve downstream utility [16].
@@
Hybrid strategies have emerged to combine complementary strengths of multiple generators and augmentations [17]. Torfi and Fox integrated correlation-capturing techniques with GANs to improve realism in healthcare records [18]. Calibration procedures, including moment matching and histogram matching, can reduce distributional drift between synthetic and real data [19]. Aligning feature-wise distributions via histogram-based adjustments is especially useful when models fit global structure but miss marginals [20]. However, existing pipelines often leave gaps that motivate a unified approach to generation and calibration for healthcare tabular data [21].
The primary research question is whether a hybrid model that integrates multiple augmentation techniques with novel calibration strategies can produce high-fidelity synthetic data while preserving downstream utility on real healthcare tasks [22]. We hypothesize that combining noise injection, interpolation, Gaussian Mixture Model (GMM) sampling, Conditional Variational Autoencoder (CVAE) sampling, and SMOTE, followed by targeted calibration, will more closely match real distributions than single-method baselines [23].
We evaluate distributional similarity using metrics such as Wasserstein distance and the Kolmogorov–Smirnov statistic [24]. We benchmark against widely used tabular synthesizers implemented in the Synthetic Data Vault ecosystem [25]. The remainder of this article is organized as follows: Section 1 reviews related work, Section 2 details the methodology, Section 3 presents experiments and analysis, and Section 4 discusses implications, limitations, and future directions.
@@
In imaging, Frid-Adar et al. demonstrated that adversarial augmentation can improve lesion classification by enriching rare phenotypes [26]. In tabular settings, synthetic data has supported fraud detection by expanding minority-class examples [27]. Market simulations have leveraged generative models to evaluate algorithmic strategies in finance [28]. Rahman et al. applied VAE-based frameworks to electronic health records for clinically meaningful synthesis [29]. Similar generative latent-variable approaches have been used for financial time series [30]. Conditional formulations increase controllability by incorporating labels and clinical context into the synthesis process [31]. Practical adoption requires awareness of computational demands and the potential for biased generations when training data are limited or skewed [32].
Mousavi et al. showed that noise injection can mitigate class imbalance and improve generalization in low-resource regimes [33]. Interpolation-based oversampling such as SMOTE provides a simple mechanism to expand minority classes in tabular data [34]. Li et al. proposed adaptive interpolation for rare disease classification problems [35]. In fraud analytics, oversampling has increased sensitivity to minority-class patterns [36]. Nonetheless, linearity assumptions may not hold in high-dimensional or nonlinear feature spaces [37]. Wang et al. used Gaussian mixture models to capture multi-modal structure and support cluster-aware sampling for synthesis [38]. These models can be challenged by sparsity and the curse of dimensionality in complex clinical tables [39].
Deville and Särndal introduced calibration estimators to align moments with real-data targets in a transparent, auditable manner [40]. Bourou et al. discussed histogram-based calibration as a practical tool to further align empirical distributions after generation [4]. Adaptive and iterative histogram strategies refine alignment by dynamically adjusting blending over successive passes [41]. Compared with image domains, calibration for tabular healthcare data remains comparatively underexplored [39].
Open-source tooling has lowered barriers to using synthetic data in clinical research [4]. Chen et al. built a cohort-level synthetic data engine demonstrating feasibility for patient-centered outcomes studies [6]. Synthetic imaging has been used to expand rare categories for model development and evaluation [42]. Yelmen et al. created artificial human genomes using generative neural networks for privacy-preserving population studies [10]. The benchmark datasets used in this work provide established testbeds for evaluating fidelity, utility, and privacy safeguards in oncology and cardiology contexts [5]. Torfi and Fox further exemplified correlation-aware hybrid generators in healthcare records [18]. Du et al. explored probabilistic–deep hybrids for clinical time series forecasting [43].
Many current methods struggle to capture higher-order dependencies essential for clinical inference in high-dimensional tables [44]. The lack of standardized evaluation protocols complicates comparisons across methods and datasets [45]. Calibration methods tailored to tabular healthcare data remain less mature than their imaging counterparts [39]. Our framework addresses these limitations by unifying multi-technique generation with a family of calibration procedures and by evaluating fidelity, utility, and privacy on oncology and cardiology benchmarks without accessing identifiable patient data.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Data-driven healthcare depends on access to high-quality, representative datasets for building and validating robust AI models [1]. In practice, U.S. and European privacy regulations impose strict limits on the use, sharing, and secondary processing of personal health information, which constrains access to real-world data for model development [2, 3]. As a mitigation, Bourou et al. [4] highlight synthetic data generation as a practical pathway that can mirror key statistics of real datasets while reducing disclosure risk. In this study, we focus on oncology and cardiology contexts using the Breast Cancer Wisconsin (Original) and Breast Cancer Wisconsin (Diagnostic) datasets from the UCI Machine Learning Repository, together with a heart-disease cohort from the Department of Cardiology, Khulna Medical College [5].
Chen et al. and related efforts show that synthetic data can enable analyses and workflow simulations under data-access constraints while protecting identities [6] [7]. In discrete EHR settings, Choi et al. introduced medGAN to synthesize multi-label records for predictive modeling in scarce-data regimes [8]. In medical imaging, Frid-Adar et al. demonstrated that GAN-based augmentation improves diagnostic performance with limited samples[9]. Beyond these exemplars, Yelmen et al. illustrate privacy-preserving potential in genomics, underscoring the broader promise of synthetic data for ethical, regulation-aware AI in biomedicine[10].
@@
Still, as emphasized by prior work [11], generating high-fidelity tabular healthcare data that preserves complex dependencies is challenging. Classical probabilistic models often struggle with high dimensionality and nonlinearity [12]. While GAN-based approaches advance realism [13] and variational autoencoders provide a complementary probabilistic route [14], practical deployment can be hindered by mode collapse, overfitting, and limited training data [15]. Consequently, many pipelines rely on post-generation calibration to retain downstream utility [16].
@@
Hybrid strategies have emerged to combine complementary strengths across generators and augmentations [17]. In healthcare records, Torfi and Fox integrated correlation-capturing techniques with GANs to improve realism [18]. For post-hoc correction, Deville and Särndal introduced calibration estimators to align sample moments with real-data targets [40], and Bourou et al. discussed histogram-based calibration as a practical tool to reduce distributional drift [4]; related work formalizes moment and histogram matching for robust alignment [19, 20]. Despite this progress, gaps often persist between generation and calibration stages, motivating unified workflows tailored to healthcare tabular data [21].
@@
Evidence across domains reinforces a hybrid-and-calibrate paradigm. In imaging, Frid-Adar et al. enriched rare phenotypes via adversarial augmentation to improve lesion classification [26]. In tabular settings, studies have applied synthetic data to bolster fraud detection through minority-class expansion [27] and have leveraged generative models for market simulations in finance [28]. Within healthcare records, Rahman et al. employed VAE-based frameworks to produce clinically meaningful cohorts, while latent-variable formulations have also been explored for financial time series [29][30]. Conditional models further increase controllability by incorporating labels and clinical context [31]. At the same time, practical adoption must account for computational demands and the risk of biased generations under limited or skewed training data [32].
Among augmentation tools, Mousavi et al. showed that noise injection can mitigate class imbalance and improve generalization in low-resource regimes [33]. Interpolation-based oversampling such as SMOTE remains a simple, effective mechanism for expanding minority classes [34]; Li et al. proposed adaptive interpolation for rare-disease classification [35], and subsequent work reported gains in fraud analytics [36], though linearity assumptions can break down in high-dimensional or nonlinear spaces [37]. To capture multi-modality, Wang et al. used Gaussian Mixture Models (GMMs) for cluster-aware sampling [38], while noting challenges from sparsity and the curse of dimensionality in complex clinical tables [39]. On the calibration front, adaptive and iterative histogram strategies further refine alignment through dynamic, successive adjustments [41]. Compared with imaging, calibration methods purpose-built for healthcare tabular data remain comparatively underexplored [39]. Open-source tooling has lowered barriers to adoption [4], cohort-level engines have demonstrated feasibility for patient-centered outcomes research [6], synthetic imaging has expanded rare categories for model development [42], and probabilistic–deep hybrids for clinical time series (e.g., Du et al.) provide additional evidence for combining model families [43]. The benchmark datasets employed here serve as established testbeds for assessing fidelity, utility, and privacy safeguards in oncology and cardiology [5].
Even so, many current methods struggle to capture higher-order dependencies essential for clinical inference in high-dimensional tables [44], and the absence of standardized evaluation protocols complicates fair comparisons across methods and datasets [45]. Motivated by these gaps, we ask whether a unified hybrid framework that integrates multiple augmentation techniques with targeted calibration can produce high-fidelity synthetic data while preserving downstream utility on real healthcare tasks [22]. We hypothesize that combining noise injection, interpolation, GMM sampling, Conditional Variational Autoencoder (CVAE) sampling, and SMOTE followed by moment- and histogram-based calibration will more closely match real distributions than single-method baselines [23]. We evaluate distributional similarity using Wasserstein distance and the Kolmogorov–Smirnov statistic [24] and benchmark against widely used tabular synthesizers in the Synthetic Data Vault ecosystem [25]. The remainder of this article is organized as follows: Section 1 reviews related work, Section 2 details the methodology, Section 3 presents experiments and analysis, and Section 4 discusses implications, limitations, and future directions.